# Distribution of square-free palindromes1

### 03 Mar 2026 v1

@@ Abstract.
$\operatorname{c}_{q}(n)$ is Ramanujan’s sum
$\Omega(n)$ is the number of prime factors of $n$, counted with multiplicity.
@@ 1. Introduction
Palindromes have been intertwined with human culture for over two millennia [^2] Reversible wordplay and verses designed to be read in more than one direction were already common in antiquity; cf. Martial [41] and the discussion of related traditions in Kwapisz [36]. For early inscriptional evidence, the ROTAS/SATOR word-square is attested at Pompeii (pre-62 CE); see Encyclopedia Britannica [24].. In recent decades, they have attracted considerable attention in mathematics [4, 5, 6, 7, 10, 11, 12, 13, 15, 33, 35, 39, 50, 56]. Although it remains unknown if there are infinitely many $b$-palindromic primes, there has been progress on both the additive and multiplicative sides.
On the additive side, Banks [4] initiated the study of representing integers as sums of palindromes, and Cilleruelo-Luca-Baxter [12] proved that every natural number is a sum of at most three $b$-palindromes when $b\geq 5$; see also Rajasekaran-Shallit-Smith [50] for the remaining small bases. On the multiplicative side, Col [15] proved there exists an integer $k_{b}$, depending only on $b$, such that infinitely many $b$-palindromes $n$ satisfy $\Omega(n)\leq k_{b}$. This was the first result of its kind. More recently, it was shown in [56] that one may take $k_{b}=6$ for every base $b$, and that $b$-palindromes have level of distribution $1/5$, uniformly in $b$.
Johnston-Kerr [35] recently derived asymptotics, under a standard coprimality constraint, for the number of square-free $b$-palindromic numbers up to $x$, thus answering a question of Banks-Shparlinski [6]. Their argument adapts the setup of Cilleruelo-Luca-Shparlinski [13] and then applies van der Corput-type methods to control certain critical ranges of square-divisor sizes. For readers unfamiliar with the van der Corput method (including the $q$-van der Corput variant) see for instance Iwaniec-Kowalski [34].
@@ 1.1. Main results
Theorem 1.1 follows from the content in Sections 3, 4, 5, 6, 7, and relies heavily on exponential-sum estimates. Some of our methodology is detailed in the upcoming subsections.
@@
Theorem 1.2 follows from Theorem 1.1 and Proposition 8.1. The latter relies on the $L^{1}$ and moment bounds in [56] as well as on the estimate of Baier-Zhao [1] (see also Baker [2]) for the large sieve with square moduli. We have not exploited the averaging over $q$ sufficiently and the content in Sections 4, 5, 6, 7 and some of Section 8 can be adapted to this end.
@@ 1.2. Structural ingredients
$d_{j}(n)=d_{\lfloor\log_{b}n\rfloor-j}(n)\qquad\forall\ \ 0\leq j<\lambda$
@@ 3. Preparatory lemmas
For any (possibly negative) integers $c,d,k,\ell,q$ with $q\geq 1$ and $k\neq\ell$,
@@
Follows directly from Lemmas 3.1, 3.6, 3.7, 3.4. ∎
@@ 4. Square divisors and quasi-palindromes
for any fixed $\epsilon>0$ and uniformly in $\operatorname{\mathfrak{a}}$, where
@@ 6. Large conductors relative to NN
The last inequality holds since $r\ll qX^{\epsilon}$ by assumption. Then by (101) and (102), we are thus in the situation of Lemma 3.12 with $X^{\delta}\ll T\ll sX^{\epsilon}$, $\|\eta^{(j)}\|_{\infty}\ll_{j}1$ for each $j\geq 0$ (by assumption) and $r=1$. Conjugating if necessary so that $T>0$, Lemmas 3.12, 3.7 and 3.4 imply
@@ 8. Medium and small square divisors
$\ll_{\epsilon}\left(qD^{3}+Kb^{2N}\sqrt{D}\right)b^{2(K-1)N}\left(1+O\left(\dfrac{1}{\sqrt{K}}+\dfrac{b^{2}}{K}\right)\right)^{2N}\left(b^{N}DqK\right)^{\epsilon}$
@@
$\sum_{d\leq D}\sum_{h(qd^{2})}^{*}\Phi_{N}^{2K}\left(\dfrac{h}{qd^{2}}+\beta\right)\ll_{\epsilon}\left(b^{N}DqK\right)^{\epsilon}\left(qD^{3}+Kb^{2N}\sqrt{D}\right)\sum_{0\leq n\leq Kb^{2N}}|\gamma_{n}|^{2}.$
By Parseval’s identity and the 11-periodicity of $\Phi_{N}$,
@@
With regards to $\Sigma_{2}$, Lemmas 8.7 and 3.16 imply
@@
$E(Q,D)=\sum_{\begin{subarray}{c}d\sim D\\ q\sim Q\\ (dq,b^{3}-b)=1\\ (d,q)=1\end{subarray}}\sup_{y\leq x}\max_{\begin{subarray}{c}(a,q)=1\\ d^{2}\mid a\end{subarray}}\left|\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(y)\\ n\equiv a(qd^{2})\end{subarray}}1-\dfrac{|\mathscr{P}_{b}^{*}(y)|}{qd^{2}}\right|.$
@@ 10. Acknowledgements
We thank Universidad de Cantabria for its warm hospitality and for making our time in Santander so memorable, both personally and mathematically. We are especially grateful to Ana Gómez-Pérez, Domingo Gómez-Pérez and Isabel Pirsic for their exceptional kindness. We thank Igor E. Shparlinski for bringing the problem to our attention and for reading an earlier draft of this work. We thank Daniel Panario for earlier discussions and reading a preliminary draft. We thank Daniel R. Johnston and Bryce Kerr for collegiate discussions and encouragement.

### 30 Mar 2026 v2

@@ 1. Introduction
Palindromes have been intertwined with human culture for over two millennia [^2] Reversible wordplay and verses designed to be read in more than one direction were already common in antiquity; cf. Martial [41] and the discussion of related traditions in Kwapisz [36]. For early inscriptional evidence, the ROTAS/SATOR word-square is attested at Pompeii (pre-62 CE); see Encyclopedia Britannica [24].. In recent decades, they have attracted attention in mathematics [4, 5, 6, 7, 10, 11, 12, 13, 15, 33, 35, 39, 50, 56]. Although it remains unknown if there are infinitely many $b$-palindromic primes, there has been progress on both the additive and multiplicative fronts.
On the additive side, Banks [4] initiated the study of representing integers as sums of palindromes, and Cilleruelo-Luca-Baxter [12] proved that every natural number is a sum of at most three $b$-palindromes when $b\geq 5$; see also Rajasekaran-Shallit-Smith [50] for the remaining small bases. On the multiplicative side, Col [15] proved there exists an integer $k_{b}$, depending only on $b$, such that infinitely many $b$-palindromes $n$ have at most $k_{b}$ prime factors. This was the first result of its kind. More recently, it was shown in [56] that one may take $k_{b}=6$ for every base $b$, and that $b$-palindromes have level of distribution $1/5$, uniformly in $b$.
Johnston-Kerr [35] recently derived asymptotics, under a standard coprimality constraint, for the number of square-free $b$-palindromes up to $x$, thus answering a question of Banks-Shparlinski [6]. Their argument adapts the setup of Cilleruelo-Luca-Shparlinski [13] and then applies van der Corput-type methods to control certain critical ranges of square-divisor sizes. For readers unfamiliar with the van der Corput method (including the $q$-van der Corput variant) see for instance Iwaniec-Kowalski [34].
@@ 1.1. Main results
Theorem 1.1 follows from Propositions 4.1, 5.1, 6.1, relying heavily on exponential-sum estimates. Some of our methodology is detailed in the upcoming subsections.
The case when $q=1$ above and $Q=1$ below was considered recently by Johnston-Kerr [35] who showed that, for any $N\geq 1$, the left hand side of (4) is $\ll_{b,\epsilon}|\Pi_{b}(L)|^{1+\epsilon}/N^{\epsilon_{0}}$ for some $\epsilon_{0}>0$ (implicit from their work after modifications to Proposition 10.1 in [56]).
@@
Theorem 1.2 follows from Theorem 1.1 and Proposition 8.1. The latter relies on the machinery for $6$-almost-primes in [56] as well as on the estimate of Baier-Zhao [1] (see also Baker [2]) for the large sieve with square moduli.
Theorem 1.1 and Proposition 8.1, although powered by rather different tools and arguments, converge to the exponent $1/16$ of Theorem 1.2 as follows. First it can be shown that Theorem 1.2 holds if we have
(7) $\sum_{\begin{subarray}{c}q\sim Q\\ (q,m_{b})=1\end{subarray}}\sup_{\begin{subarray}{c}(a,q)=1\\ y\leq x\end{subarray}}\sum_{\begin{subarray}{c}d\sim D\\ (d,m_{b}q)=1\end{subarray}}\left|\sum_{n\in\mathscr{P}_{b}^{*}(y)}\left(\mathbf{1}_{d^{2}\mid n\equiv a(q)}-\dfrac{1}{qd^{2}}\right)\right|\ll_{b,\epsilon}\sqrt{x}\exp\left(-\sigma^{\prime}\sqrt{\log x}\right)$
for any $Q\leq x^{1/16-\epsilon}$ and $D\leq\sqrt{x}$. This follows from $\mu^{2}(n)=\sum_{d^{2}\mid n}\mu(d)$, an Euler product expansion of $\sum_{\begin{subarray}{c}(d,m_{b}q)=1\end{subarray}}\mu(d)/d^{2}$ and dyadic decompositions.
  • Theorem 1.1 implies that if $DQ\gg x^{1/4-\epsilon}$, then (7) holds if $Q\ll x^{1/16-\epsilon}$.
  • Proposition 8.1 implies that if $DQ\ll x^{1/4-\epsilon}$, then (7) holds if $Q\ll x^{1/15-\epsilon}$.
@@ 1.2. Structural ingredients
$d_{j}(n)=d_{\lfloor\log_{b}n\rfloor-j}(n)\qquad\forall\ \ 0\leq j<\lambda,$
@@ 3. Preparatory lemmas
Lemma 3.6 (Multiplicative property).
For any integers $c,d,q,r$ with $q,r\geq 1$ and $(q,r)=1$,
$\operatorname{K}_{2}(c,d;qr)=\operatorname{K}_{2}(c\overline{r},d\overline{r};q)\operatorname{K}_{2}(c\overline{q},d\overline{q};r).$
Proof.
Follows from the Chinese remainder theorem. ∎
@@
For any (possibly negative) integers $c,d,k,\ell,q$ with $q\geq 1$, $k\neq\ell$ and $k,\ell\neq 0$,
@@
Follows directly from Lemmas 3.1, 3.7, 3.8, 3.4. ∎
@@ 4. Square divisors and quasi-palindromes
for any fixed $\epsilon>0$, where
@@ 6. Large conductors relative to NN
The last inequality holds since $r\ll qX^{\epsilon}$ by assumption. Then by (102) and (103), we are thus in the situation of Lemma 3.13 with $X^{\delta}\ll T\ll sX^{\epsilon}$, $\|\eta^{(j)}\|_{\infty}\ll_{j}1$ for each $j\geq 0$ (by assumption) and $r=1$. Conjugating if necessary so that $T>0$, Lemmas 3.13, 3.8 and 3.4 imply
@@ 8. Medium and small square divisors
$\ll_{\epsilon}\left(qD^{3}+Kb^{2N}\sqrt{D}\right)b^{2(K-1)N}\left(1+\dfrac{q}{Kb^{2N}}\right)\left(1+O\left(\dfrac{1}{\sqrt{K}}+\dfrac{b^{2}}{K}\right)\right)^{2N}\left(b^{N}DK\right)^{\epsilon}$
@@
$\sum_{d\leq D}\sum_{h(qd^{2})}^{*}\Phi_{N}^{2K}\left(\dfrac{h}{qd^{2}}+\beta\right)$
$\ll_{\epsilon}\left(b^{N}DK\right)^{\epsilon}\left(qD^{3}+Kb^{2N}\sqrt{D}\right)\left(1+\dfrac{q}{Kb^{2N}}\right)\sum_{0\leq n\leq Kb^{2N}}|\gamma_{n}|^{2}$
for any $\epsilon>0$. By Parseval’s identity and the 11-periodicity of $\Phi_{N}$,
@@
With regards to $\Sigma_{2}$, Lemmas 8.7 and 3.17 imply
@@
$E(Q,D)\leq\sum_{\begin{subarray}{c}d\sim D\\ q\sim Q\\ (dq,b^{3}-b)=1\\ (d,q)=1\end{subarray}}\sup_{y\leq x}\max_{\begin{subarray}{c}(a,q)=1\\ d^{2}\mid a\end{subarray}}\left|\sum_{\begin{subarray}{c}n\in\mathscr{P}_{b}^{*}(y)\\ n\equiv a(qd^{2})\end{subarray}}1-\dfrac{|\mathscr{P}_{b}^{*}(y)|}{qd^{2}}\right|.$
@@ 10. Acknowledgements
We thank Universidad de Cantabria for its warm hospitality and for making our time in Santander so memorable, both personally and mathematically. We are especially grateful to Ana Gómez-Pérez, Domingo Gómez-Pérez and Isabel Pirsic for their exceptional kindness. We thank Igor E. Shparlinski for bringing the problem to our attention and for reading an earlier draft of this work. We thank Daniel Panario for earlier discussions and reading a preliminary draft. We thank Daniel R. Johnston and Bryce Kerr for collegiate discussions and encouragement. Part of the work was supported by Cátedra Universidad de Cantabria - INCIBE de nuevos retos en ciberseguridad, financed by European Union NextGeneration-EU, the Recovery Plan, Transformation and Resilience, through INCIBE. A preliminary stage of this project was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC).